The sum of two numbers is 94. The second is 6 more then 3 times the first. What are the two numbers.
Answer by checkley75(3666)

X+Y=94, X=94-Y
Y=6+3X NOW REPLACE X BY (94-Y) IN THIS EQUATION & SOLVE FOR Y.
Y=6+3(94-Y)
Y=6+282-3Y
Y+3Y=288
4Y=288
Y=288/4
Y=72 ANSWER FOR THE LARGER NUMBER.
X=94-72
X=22 ANSWER FOR THE SMALLER NUMBER.
PROOF
72=6+3*22
72=6+66
72=72